# GENIAL: Generative Design Space Exploration via Network Inversion for Low Power Algorithmic Logic Units

### v1

@@ III-C4 Switching Activity Model
  3. 3.
    Transition counting: count how many times each net state flips between consecutive cycles. Multiply each flip count with the total transistor count of that net’s fan-out cells — approximating the real hardware cost of a individual wires transitions. This is illustrated Fig. 4.
  4. 4.
@@ III-D1 Quality of Results Prediction
To recommend new designs, it is necessary to predict their QoR. To do so, we train a deep learning model to perform a regression on the design scores. The model takes as input the 2D tensor representation of a design’s encoding and outputs a scalar value, which corresponds to the predicted score for that encoding. Each known design is assigned a score, which is the normalized value of a target QoR metric. The encoding-score pairs serve as supervised training samples.
III-D2 Model Architecture
Fig. 5 depicts the architecture of the surrogate model, which consists in a PointNet-based[28] embedding followed by a stack of Transformer encoder layers[38].
Embedding: The PointNet, a sequence of one-dimensional convolution layers (C1D), converts each individual binary representation of the input encoding into a fixed size vector—also called a token. A learned vector of the same dimension which serves as a global representation of the design for the transformer model—commonly called [CLS] token—is prepended to this sequence of vectors. Finally, a learned sequence of absolute positional encoding is added element-wise to the token sequence.
Transformer and Head: the embedded sequence of tokens is then fed into a stack of Transformer encoder blocks. Each block performs non-masked multi-head self-attention and outputs a new sequence of tokens. At the output of the stack, only the vector corresponding to the [CLS] token of the final output sequence is passed through a series of linear layers, the model head, which outputs a single scalar value: the predicted score.
@@ III-D3 Training Pipeline
Training a model from random weights has been proven difficult as the model needs to learn to efficiently represent each input vector in the embedding space, and then learn to predict QoR metrics out of this latent space. To accelerate convergence and improve final model accuracy, we design a two-step training pipeline: 1) Self-Supervised Learning (SSL) to initialize meaningful embedding and attention weight spaces, and 2) finetuning on the real task.
Self-Supervised Pretraining: The SSL pretraining phase consists in a simple permutation prediction task. It involves predicting the positional indices of the permuted binary vectors. The TC encoding is used as the template encoding, with indices ordered from 0 to 15. For each training sample, a random permutation of the 16 binary representations and associated indices is performed. The permuted sample is passed through the Embedding and Transformer blocks. A 2-layer perceptron (MLP) receives the representations and predicts the positional indices corresponding to the permutation of the binary encodings. The target is represented as a one-hot encoded sequence of length 16, and a cross entropy loss is applied. Once the pretraining is completed, the Embedding and Transformer weights are retained to initialize the model for finetuning, and the MLP head is discarded.
Finetuning: In this phase, the model is trained on the real task of predicting QoR metrics. The target score is derived the normalized QoR metric of each known design, and the Mean Square Error (MSE) is used as the loss function. We use standardization to normalize the QoR values. Via backpropagation and weight optimization with the Adam optimizer, the model then learns to map an encoding input (represented as a 2D tensor) with a score.
Data Augmentation: We exploit some notable properties of digital designs to derive a data augmentation scheme. 1) Swapping the positions of a logic unit’s input bits leaves its fundamental function unchanged. Similarly, permuting the columns of the encoding tensor does not change the logic function of the associated design. There are 24 possible permutations of 4 columns. 2) Adding inverters in front of all input bits is essentially free in terms of added logic gates. Furthermore, the SwAct depends on signal toggles, and is independent of their polarity. Hence, negating the encoding (by exchanging 0s and 1s) should likewise preserve the implementation cost.
These two transformations preserve the legality of an encoding. Thus each known design can be augmented with 2×24=482$\times$ 24=482 × 24 = 48 different encodings resulting in the same QoR.
@@ III-E Design Recommendation
The foundation of our framework relies on training a surrogate model able to predict a score that represents the quality of a design from its abstracted representation.
@@ III-E1 Naive Approach
The Naive Approach (NA) is an intuitive baseline that involves: 1) randomly generating a large set of candidate encodings, 2) predicting their score using the surrogate model, and 3) selecting and processing only the top designs based on predicted scores. This process is repeated iteratively over several generations: at each round, newly generated designs are added to the existing dataset, and the surrogate model is finetuned with the updated data. While this iterative refinement is expected to progressively guide the search toward higher quality design (QoR), in practice, the performance of this approach remains significantly worse than the Model Inversion (MI) approach described next.
III-E2 Model Inversion
The MI method follows a gradient-based optimization approach. It involves: 1) initializing the input tensors with noise, 2) performing a forward pass through the model, 3) backpropagating gradients to the inputs, and 4) updating the input to minimize the loss. This procedure is illustrated Fig. 6. During optimization, the model weights remain frozen. Only the input tensor, representing the design encoding, is updated by the Adam optimizer. The primary component of the loss is the model’s output itself, so the optimizer attempts to minimize this output by updating the input. To generate diverse design candidates, the initial input tensors are filled with a constant value of 0.5 and perturbed with small Gaussian noise. This allows for multiple unique encodings to be generated while using the same fixed model weights.
An input encoding tensor is considered legal only if it contains every possible binary representation exactly once. To increase the likelihood of generating such legal encodings, we introduce two regularization terms: 1) an ”attraction” loss, which pulls each input vector toward all possible binary target vectors, encouraging convergence to binary values (0s and 1s), and 2) a repulsion loss, which pushes input vectors away from one another, discouraging vectors from converging to the same target vector. These regularization terms are weighted by a scalar coefficient that balances the trade-off between minimizing the primary score loss and enforcing the attraction-repulsion constraints. After 1500 epochs, most of the vectors of an input converge to valid binary values, though some remain in an intermediate state. A post-processing step is thus applied to legalize the encoding. Specifically, for each floating vector that has not converged to a binary target, we compute its distance to the remaining unassigned binary target vectors. The vector with the smallest distance to any available target vector is assigned to that target, and both are removed from their respective candidate pools. This process is repeated iteratively until all vectors have been assigned and the encoding is legal.
Figure 6: Model inversion setup
@@ IV-A Experimental Setup
The experiment involves the following steps. Model Pretraining: The model is pretrained using SSL for 200 epochs. Initial Generation: A set of random valid designs is generated and processed to extract the QoR, forming the initial dataset. Model Finetuning: The model is trained on the full available dataset for 200-400 epochs. The checkpoint with the lowest validation loss is selected for the recommendation. Prototype Generation: recommended candidate encodings are generated via MI for 1500 epochs. QoR Extraction: The design corresponding to the selected encodings are generated as LUTs. They are synthesized with one of the available flows, and their QoR are extracted. Iteration: The newly evaluated designs are added to the dataset, and the process is repeated for another round.
At each generation round, the dataset used for finetuning the network is expanded with additional designs, leading to two key effects: it improves the maximum attainable accuracy of the surrogate model, and it enriches the dataset with more diverse design knowledge. Consequently, surrogate-model-based recommendations generally improve over successive generation rounds.
@@ IV-B OPENR Optimization for Power
In this experiment, we run GENIAL using the OPENR flow, targeting Power (P) as the QoR metric to optimize. Results are obtained by synthesizing and simulating each design with OpenRoad Flow Scripts in the synthetic ASAP7 technology[8]. Each design is compiled and simulated with Verilator over 20k randomly generated stimuli (i.e. input operand pair switches), with a normal distribution of mean 0 and standard deviation ($sigma$) 3[3]. The values are clipped to the range (-8;7) and rounded to nearest integer. Simulation waveforms are then used to perform power extraction. Fig. 7 illustrates the evolution of the minimum and mean power (P) across generation rounds, showing that GENIAL consistently reduces power consumption over iterations. However, the LUT-based SME and TC encodings were not optimized as effectively as their hand-crafted counterparts implemented with Verilog arithmetic operators, as shown in Table II. The LUT SME design exhibits higher power than the LUT TC, which should not be the case, as highlighted by power extraction results realized on handmade designs (see Table II). Since GENIAL’s ability to discover meaningful designs strongly depends on the accuracy of the QoR metrics used to train the QMP, higher-effort LSO in the loop is necessary.
TABLE II: OPENR Synthesis and Power Extraction Results
@@ IV-C FLOWY Optimization for Switching Activity
We thus conducted a larger experiment with FLOWY in the loop. Fig. 8 presents the results. The initial dataset consists of 10k randomly generated designs, to which we add 577 designs uniformly sampled from a separate dataset obtained by previously running GENIAL with OPENR targeting NT as the QoR metric. This was done to increase dataset diversity and maximize the probability of success. To avoid biasing the search, we make sure that this initial dataset does not contain the TC or SME encodings. We applied all proposed quality enhancement techniques, including SSL pretraining, data augmentation, and even over-generation during MI and subsequent selection of the 25% best predicted encoding. For the first 10 generation rounds, we recommend 2k designs per iteration. Subsequently, as training time became more expensive than design synthesis, we increased the number of design per generation round to 4.5k. The final simulation for each best design in every chain was performed with 20k stimuli, with operand values randomly sampled from a normal distribution with $\sigma=3.0$, as above.
After generation round 13 (GR13), we observe convergence towards an encoding that exhibits significant symmetry around zero, but different from the SME. A similar encoding was found at generation round 14 (GR14). All data points are illustrated Fig. 8, and the two encodings mentioned here are depicted Fig. 2.
We find several solutions at the Pareto-front of SwAct-v-NMIG. Particularly, we find solutions expected to both be smaller and use less power than TC at the same time. Some also use 12% less area than SME for only 13% more power.
We then perform a higher effort LSO (with 15k steps per chain and 120 chains per design) on selected encodings and present the results Table III (a). They show that the obtained encoding performs nearly as well as the SME encoding and consume less power than the TC by 18%.
@@
TABLE III: FLOWY Higher Effort Synthesis Results ($\sigma=3$)
@@ V-A Metric Correlations
GENIAL is rigorously validated through a series of experiments. Since FLOWY is relatively slow (processing a single encoding takes about 15min to 20min on 18 CPUs), we evaluate our methods using the faster ESPRS or OPENR flows. We observe a strong correlation between most metrics across the different flows, as shown in Fig. 9, confirming that ESPRS and OPENR flows serve as reliable proxies for tuning our method.
Figure 10: Comparison of Naive approach vs Surrogate Model Inversion
@@ V-B1 Naive versus Model Inversion
To compare MI against the NA, we evaluate both methods under equivalent compute duration, over three trials. Since MI requires several backward passes to converge on meaningful encoding prototypes, generating 10k samples takes approximately 25 minutes. Therefore, for each generation round, the NA method was run for 25 minutes, resulting in about 10M samples. From these, only the top 10k designs based on predicted scores by the QMP were retained and processed. Results of this experiment, using the OPENR flow with transistor count (NT) as the target QoR metric, are shown in Fig. 10. After just two generation rounds (30k processed samples), MI already discovers encodings with lower minimum NT than NA. After five rounds (60k processed samples), the MI-recommended population has a lower mean NT than the naive approach, demonstrating the superior effectiveness of MI.
Figure 11: Comparison between methods
@@ V-B2 Sample Efficiency
We compare our method against three others by limiting the number processed samples, over three trials. Fig. 11 shows the expected minimum value reached as a function of the sample budget for two metrics: CPX and NT. The methods compared are: a) Random Search: uniformly sample legal encodings. b) Binary Switching Search[46]: starting from a random legal encoding, evaluate all “swap neighbors” (encodings obtained by swapping two rows in the input representation), move to the lowest‐cost neighbor (with a 25% random move chance), and restart from a new random encoding every 5000 samples processed. c) Naive Approach with Surrogate Models: as described above. d) GENIAL: using 10k initial samples and 10k samples per generation. Excepted at very low budgets (where the model has insufficient training data), GENIAL consistently finds the lowest-cost encodings, demonstrating superior sample efficiency.

### v2

@@ III-C5 Switching Activity Model
  3. iii.
    Transition counting: count how many times each net state flips between consecutive cycles. Multiply each flip count by the total transistor count of that net’s fan-out cells, thereby approximating the real energy cost of an individual wire’s transitions. This is illustrated in Fig. 4.
  4. iv.
@@
We validate SwAct as a proxy via strong correlations to Power and even Number of Transistors (Fig. 9). Its use in FLOWY enables accurate selection of final designs without the need to fully simulate all candidates.
@@ III-D1 Quality of Results Prediction
To recommend new designs, it is necessary to predict their QoR. Accordingly, we train a deep learning network to perform a regression on the design scores. The network takes as input the 2D tensor representation of a design’s encoding and outputs a scalar value, which corresponds to the predicted score for that encoding. Each known design is assigned a score, which is the standardized value of a target QoR metric. The encoding-score pairs serve as supervised training samples.
III-D2 Network Architecture
Fig. 5 depicts the architecture of the surrogate model, which consists in a PointNet-based[Qi_2017_CVPR] embedding followed by a stack of Transformer encoder layers[NIPS2017_3f5ee243].
Embedding: The PointNet, a sequence of one-dimensional convolution layers (C1D), converts each individual binary representation of the input encoding into a fixed-size vector, also called a token. A learned vector of the same dimension, which serves as a global representation of the design for the Transformer model and commonly called the [CLS] token, is prepended to this sequence of vectors. Finally, a learned sequence of absolute positional encodings is added element-wise to the token sequence.
Transformer and Head: The embedded token sequence is subsequently fed into a stack of Transformer encoder blocks. Each block performs non-masked multi-head self-attention and outputs a new sequence of tokens. At the output of the stack, only the vector corresponding to the [CLS] token of the final output sequence is passed through a series of linear layers, the model head, which outputs a single scalar value: the predicted score.
@@ III-D3 Training Pipeline
Training a network from random initialization is challenging, as it must learn both to efficiently represent each input vector in the embedding space and to predict QoR metrics from the resulting latent representation. To accelerate convergence and improve final accuracy, we adopt a two-stage training pipeline, as well as a data augmentation scheme.
Self-Supervised Pretraining: The Self-Supervised Learning (SSL) pretraining phase consists of a simple permutation prediction task. It involves predicting the positional indices of the permuted binary vectors. The TC encoding is used as the template encoding, with indices ordered from 0 to 15. For each training sample, a random permutation of the 16 binary representations and associated indices is performed. The permuted sample is passed through the Embedding and Transformer blocks. A 2-layer perceptron (MLP) receives the [CLS] token output representation and predicts the positional indices corresponding to the permutation of the binary encodings. The target is represented as a one-hot encoded sequence of length 16, and a cross entropy loss is applied. Once the SSL pretraining task is completed, the Embedding and Transformer weights are retained to initialize the model for finetuning, and the MLP head is discarded.
Finetuning: In this phase, the model is trained on the real task of predicting QoR metrics. The target score corresponds to the standardized QoR value of each known design, and the Mean Square Error (MSE) serves as the loss function. The network weights are optimized via backpropagation with the Adam optimizer [kingma2017adammethodstochasticoptimization], and the model learns to map an encoding input (represented as a 2D tensor) to its corresponding score.
Data Augmentation: We exploit some notable properties of digital designs to derive a data augmentation scheme:
  1. i.
    Swapping the positions of a logic unit’s input bits leaves its fundamental function unchanged. Similarly, permuting the columns of the encoding tensor does not change the logic function of the associated design. There are 24 possible permutations of 4 columns.
  2. ii.
    Adding inverters in front of all input bits is essentially free in terms of added logic gates. Furthermore, the SwAct depends on signal toggles, and is independent of their polarity. Hence, negating the encoding (by exchanging 0s and 1s) should likewise preserve the implementation cost.
These two transformations preserve the legality of an encoding. Thus each known design can be augmented with $2\times 24=48$ different encodings resulting in the same QoR.
@@ III-E Design Recommendation
With the surrogate model finetuned, we leverage its predictions to guide the search toward promising designs and compare four different strategies for generating new candidates.
@@ III-E1 Naive Approach
The Naive Approach (NA) is an intuitive baseline that involves:
  1. i.
    randomly generating a large set of candidate encodings,
  2. ii.
    predicting their score using the surrogate model, and
  3. iii.
    selecting and processing only the top designs based on predicted scores.
This process is repeated iteratively over several generations: at each round, newly generated designs are added to the existing dataset, and the surrogate model is finetuned with the updated data. While this iterative refinement is expected to progressively guide the search toward higher quality designs (QoR), in practice, the performance of this approach remains significantly worse than the Network Inversion (NI) approach described next.
III-E2 Network Inversion
The NI method employs a gradient-based optimization approach. As illustrated in Fig. 6, it proceeds by:
  1. i.
    initializing the input tensors with noise,
  2. ii.
    performing a forward pass through the network,
  3. iii.
    backpropagating gradients to the inputs, and
  4. iv.
    updating the inputs to minimize the loss.
During optimization, the network weights remain frozen; only the input tensor, representing the design encoding, is updated using the Adam [kingma2017adammethodstochasticoptimization] optimizer. The primary component of the loss is the surrogate model’s output itself, since the objective is to identify input encodings that yield low QoR values. The optimizer iteratively adjusts the input representation to reduce the loss, decreasing the surrogate prediction value. To ensure that distinct designs are generated, each input tensor is initialized with a constant value of 0.5 and perturbed with small Gaussian noise. This initialization strategy enables the generation of multiple distinct encodings while using the same fixed model parameters.
An input encoding tensor is considered legal if it contains every possible binary representation exactly once. To increase the likelihood of generating such legal encodings, we introduce two regularization terms:
  1. i.
    An attraction loss, which pulls each input vector toward all possible binary target vectors, encouraging convergence to binary values (0s and 1s).
  2. ii.
    A repulsion loss, which pushes input vectors away from one another, discouraging vectors from converging to the same target.
These regularization terms are weighted by a scalar coefficient that balances the trade-off between minimizing the QoR loss and enforcing the attraction-repulsion constraints. After 1500 optimization epochs, most of the vectors of an input converge to valid binary values, though some remain in intermediate states. A post-processing step is thus applied to legalize the encoding. Specifically, for each non-binary vector, we compute its distance to the remaining unassigned binary target vectors. The vector-target pair with the smallest distance is then matched, and both are removed from their respective candidate pools. This process is repeated iteratively until all vectors are assigned, yielding a fully legal encoding.
Figure 6: Network inversion setup
@@ IV-A Experimental Setup
The experiment involves the following steps.
Model Pretraining: The model is pretrained using SSL for 200 epochs.
Initial Generation: A set of random valid designs is generated and processed to extract the QoR, forming the initial dataset.
Model Finetuning: The model is trained on the full available dataset for 200-400 epochs. The checkpoint with the lowest validation loss is selected for the recommendation.
Prototype Generation: Recommended candidate encodings are generated via NI for 1500 epochs. These 1500 NI epochs yield  99% vectors binary-converged before legalization; beyond that we saw diminishing returns.
QoR Extraction: The designs corresponding to the selected encodings are generated as LUTs. They are synthesized with one of the available flows, and their QoR are extracted.
Iteration: The newly evaluated designs are added to the dataset, and the process is repeated for another round. At each generation round, the dataset used for finetuning the network is expanded with additional designs, leading to two key effects: it improves the maximum attainable accuracy of the surrogate model, and it enriches the dataset with more diverse design knowledge. Consequently, surrogate-model-based recommendations generally improve over successive generation rounds.
Design Methodology Reference: For clarity, we briefly restate the role of each component in GENIAL:
  1. i.
    the Design Generator converts candidate encodings into synthesizable RTL (Section III-B);
  2. ii.
    the EDA Task Launcher evaluates each RTL design using one of the three flows described in Section III-C;
  3. iii.
    the Quality Metric Predictor predicts the QoR of new designs from their encoded representation (Section III-D);
  4. iv.
    the Design Recommender uses Network Inversion to propose new encodings based on surrogate gradients (Section III-E).
@@ IV-B OPENR Optimization for Power
In this experiment, we run GENIAL using the OPENR flow, targeting Power (P) as the QoR metric to optimize. Results are obtained by synthesizing and simulating each design with OpenROAD Flow Scripts in the synthetic ASAP7 technology [CLARK2016105]. Each design is compiled and simulated with Verilator over 20k randomly generated stimuli (i.e. input operand pair switches), with a normal distribution of mean 0 and standard deviation ($\sigma=3$). The values are clipped to the range (-8, 7) and rounded to the nearest integer. Simulation waveforms are subsequently used to perform power extraction.
Fig. 7 illustrates the evolution of the minimum and mean power (P) across generation rounds, showing that GENIAL consistently reduces power consumption over iterations. However, the LUT-based SME and TC encodings were not optimized as effectively as their hand-crafted counterparts implemented with Verilog arithmetic operators, as shown in Table III. The LUT SME design exhibits higher power than the LUT TC, which should not be the case, as highlighted by power extraction results realized on handmade designs (see Table III). Since GENIAL’s ability to discover meaningful designs strongly depends on the accuracy of the QoR metrics used to train the QMP, a higher-effort LSO in the loop is necessary.
TABLE III: OPENR Synthesis and Power Extraction Results
@@
Note on LUT vs. handmade RTL. As seen in Table III (OPENR), LUT-SME exhibits higher P than LUT-TC, whereas handmade operator-based RTL shows the opposite trend. We therefore treat LUT-RTL power as a screening signal only and rely on FLOWY (high-effort LSO) and SwAct for final conclusions in Sec. IV-C.
@@ IV-C FLOWY Optimization for Switching Activity
We conduct a larger experiment with FLOWY in the loop, and the results are shown in Fig. 8. The initial dataset consists of 10k randomly generated designs, augmented with 577 additional designs uniformly sampled from a separate dataset previously generated by running GENIAL with OPENR targeting NT as the QoR metric. This augmentation increases dataset diversity and maximizes the likelihood of successful optimization. To avoid biasing the search, the initial dataset excludes both the TC or SME encodings. We apply all proposed quality-enhancement techniques, including SSL pretraining, data augmentation, and over-generation during NI, followed by a selection of the top 25% predicted encodings. For the first 10 generation rounds, 2k designs are recommended per iteration. Later, as training time became the limiting factor relative to design synthesis, we increased the number of designs per round to 4.5k. The final simulation for each best design in every chain was performed with 20k stimuli, with operand values randomly sampled from a normal distribution with $\sigma=3.0$, as above.
After generation round 13 (GR13), we observe convergence towards an encoding that exhibits significant symmetry around zero, which is different from the SME. A similar encoding was found at generation round 14 (GR14). When referring to GR13/GR14 encodings in Fig. 8, we visualize them in Fig. 2; correlations among proxies are in Fig. 9.
We find several solutions at the Pareto-front of SwAct vs NMIG. Particularly, we find solutions expected to be both smaller and to use less power than TC. Some encodings also use 12% less area than SME for only 13% more power.
We then perform a higher effort LSO (with 15k steps per chain and 120 chains per design) on selected encodings and present the results in Table IV (a). They show that the obtained encoding performs nearly as well as the SME encoding and consumes 18% less power than the TC.
@@
TABLE IV: FLOWY Higher Effort Synthesis Results ($\sigma=3$)
@@
Note that the current code implementation (available at https://github.com/huawei-csl/GENIAL) is built around Docker containers and supports multi-node, multi-CPU parallelism. This allows the search to be conducted at scale within a reasonable time frame, even when using the high-effort FLOWY synthesis (approximately one week for the full run shown in Fig. 8).
@@ V-A Metric Correlations
GENIAL is rigorously validated through a series of experiments. Since FLOWY is relatively slow (processing a single encoding takes about 15-20 min on 18 CPUs), we evaluate our methods using the faster ESPRS or OPENR flows. We observe a strong correlation between most metrics across the different flows, as shown in Fig. 9, confirming that ESPRS and OPENR flows are reliable proxies for tuning our method.
Figure 10: Comparison of Naive approach and Surrogate Model Inversion
@@ V-B1 Naive versus Model Inversion
To compare NI against the NA, we evaluate both methods under equivalent compute duration, over three trials. Since NI requires several backward passes to converge on meaningful encoding prototypes, generating 10k samples takes approximately 25 minutes. Therefore, for each generation round, the NA method was run for 25 minutes, resulting in about 10M samples. From these, only the top 10k designs based on predicted scores by the QMP were retained and processed. Results of this experiment, using the OPENR flow with NT as the target QoR metric, are shown in Fig. 10. After just two generation rounds (30k processed samples), NI already discovers encodings with lower minimum NT than NA. After five rounds (60k processed samples), the NI-recommended population has a lower mean NT than the naive approach, demonstrating the superior effectiveness of NI.
Figure 11: Comparison between methods presented in section V-B2.
@@ V-B2 Sample Efficiency
We compare our method against three others by limiting the number of processed samples across three trials. Fig. 11 shows the expected minimum value achieved as a function of the sample budget for two metrics: CPX and NT. The methods compared are:
  1. i.
    Random Search: uniformly sample legal encodings.
  2. ii.
    Binary Switching Search[64657]: starting from a random legal encoding, evaluate all swap neighbors (encodings obtained by swapping two rows in the input representation), move to the lowest‐cost neighbor (with a 25% random move chance), and restart from a new random encoding every 5000 samples processed.
  3. iii.
    Naive Approach with Surrogate Models: as described in Section III-E.
  4. iv.
    GENIAL: using 10k initial samples and 10k samples per generation.
Except at very low budgets (where the network has insufficient samples to effectively train its parameters), GENIAL consistently finds the lowest-cost encodings, demonstrating superior sample efficiency. At an equal number of processed samples, NI consistently attains lower minima and means than NA because it exploits surrogate gradients to target promising encoding manifolds directly, rather than relying on random sampling plus predictions (Fig. 11). As the dataset grows, the QMP’s accuracy improves and NI’s sample efficiency further increases.